# Meta-Learning Driven Movable-Antenna-assisted Full-Duplex RSMA for Multi-User Communication: Performance and Optimization

### v2

@@ III Problem Formulation
$\qquad||\textbf{t}_{n_{t}}^{BS}-\textbf{t}_{\hat{n_{t}}}^{BS}||\geq DS,1\leq n_{t}\neq\hat{n_{t}}\leq N_{T},$ (31c)
$\qquad||\textbf{r}_{n_{r}}^{BS}-\textbf{r}_{\hat{n_{r}}}^{BS}||\geq DS,1\leq n_{r}\neq\hat{n_{r}}\leq N_{R},$ (31d)
@@
where $\textbf{P}=[P_{u,j},P_{U}],\forall u\in\mathcal{U},\forall j\in J$, $\textbf{W}=[\textbf{W}_{c},\textbf{W}_{p,d}],\forall d\in\mathcal{D}$, $\textbf{Z}=[\textbf{Z}_{u,j},\textbf{Z}_{U}],\forall u\in\mathcal{U},\forall j\in J$, $\textbf{t}=[\textbf{t}_{n_{t}}^{BS},\textbf{t}^{UL}],\forall n_{t}\in N_{T}$, $\textbf{r}=[\textbf{r}_{n_{r}}^{BS},\textbf{r}^{DL}],\forall n_{r}\in N_{R}$. Here, (30b) refers to the power budget constraint of BS. (30c) and (30d) refer to the power budget constraint of UL UEs. Meanwhile, (30e) ensures the QoS constraints in terms of data rate requirements in both DL and UL UEs. (31b) and (30f) limit the ranges of MA movements. (30g) and (30h) associate with the receiving beamforming vectors at the BS. (30i) represents the constraint for successful decoding of common stream by all DL UEs. Finally, (31c) and (31d) ensure that minimum inter-MA distance $DS$ at the BS for practical implementation.
It should be noted that $\mathcal{P}_{1},\mathcal{P}_{2},\mathcal{P}_{3}$ are highly nonconvex optimization problem. Specifically, the objective function, QoS constraints, receiving beamforming constraint, and minimum inter-MA distance constraints make the optimization problem highly intractable. Moreover, the optimization problem results in large-scale optimization problems with an increasing number of variables with increasing UEs. Hence, traditional optimization algorithms will suffer from high computational complexity to solve the proposed optimization problem. Thus, we propose a learning-based approach called gradient-based meta-learning, aka GML, to solve the problem with low complexity.
@@ IV Proposed Solution: Gradient-based Meta Learning (GML) For Resource Allocation
Conventional data-driven meta-learning approaches, such as model-agnostic meta-learning (MAML) [22], rely heavily on extensive offline pretraining and frequent online adaptations. However, their high energy consumption and sensitivity to variations in data distribution make them impractical for dynamic, latency-sensitive environments. To address these challenges, we employ a model-driven GML framework, as proposed in [19], which eliminates the need for pretraining and focuses on trajectory optimization rather than optimizing individual variables. This approach has been demonstrated to be effective for non-convex optimization problems [23]. The GML algorithm follows a three-layer nested cyclic optimization structure, consisting of epoch, outer, and inner iterations, which are further detailed below.
@@ IV-1 Inner iterations
[…]
@@
$\mathcal{P}_{1}:\max\sum_{d\in\mathcal{D}}R_{d}(\hat{\textbf{W}},\hat{\textbf{u}},\hat{\textbf{c}})+\sum_{u\in\mathcal{U}}R_{u}(\textbf{P}^{(i,j)},\hat{\textbf{u}},\hat{\textbf{Z}}),\>$ (38)
where $\hat{\textbf{W}},\hat{\textbf{u}},\hat{\textbf{c}},\hat{\textbf{Z}}$ represent either the initialized or updated parameters. The variable P is updated in two steps. First, we compute the sum rate based on (38) where we only optimize the target variable and keep the other variables from initialization or updated from the other network. Second, $\textbf{P}^{(i,j)}$ is fed into the neural network. Then, the output $\Delta\textbf{P}_{(i,j)}$ is added to $\textbf{P}^{(i,j)}$ to update it for the next iteration. This update can be given as follows,
$\footnotesize{P}_{u,j}^{(i+1,j)}={P}_{u,j}^{(i,j)}+\Delta{P}_{u,j}^{(i,j)},$ (39)
@@
$\footnotesize X_{1}=\frac{P_{u}^{max}}{\sum_{j\in J}P_{u,j}}P_{u,j}^{(i+1,j)},$ (43)
$\footnotesize X_{2}=\frac{P_{U}^{max}}{P_{U}}P_{U}^{(i+1,j)}.$ (44)
@@
$\mathcal{P}_{2}:\max\sum_{d\in\mathcal{D}}R_{d}({\textbf{W}^{(i,j)}},\hat{\textbf{u}},\hat{\textbf{c}})+\sum_{u\in\mathcal{U}}R_{u}\hat{(\textbf{P}},\hat{\textbf{u}},\hat{\textbf{Z}}).\>$ (45)
@@
$\footnotesize\textbf{W}^{*}=\textbf{W}^{(i,j)}+\Delta\textbf{W}^{(i,j)}.$ (46)
@@
$\footnotesize\textbf{W}^{(i+1,j)}=\sqrt{\frac{P_{BS}}{\textrm{Tr}(\textbf{h}_{d}^{H}\textbf{W}^{*}(\textbf{h}_{d}^{H}\textbf{W}^{*})^{H})}}\textbf{W}^{*}.$ (47)
Receiving Beamforming Network (RBN): In RBN, we optimize Z, while keeping other variables from its initial states or updated from the other networks. Hence, the objective and update of gradient is measured as follows,
$\mathcal{P}_{3}:\max\sum_{d\in\mathcal{D}}R_{d}({\hat{\textbf{W}}},\hat{\textbf{u}},\hat{\textbf{c}})+\sum_{u\in\mathcal{U}}R_{u}\hat{(\textbf{P}},\hat{\textbf{u}},{\textbf{Z}^{(i,j)}}),\>$ (48)
$\footnotesize\textbf{Z}^{(i+1,j)}=\textbf{Z}^{(i,j)}+\Delta\textbf{Z}^{(i,j)}.$ (49)
@@
$\mathcal{P}_{4}:\max\sum_{d\in\mathcal{D}}R_{d}({\hat{\textbf{W}}},\hat{\textbf{u}},\textbf{c}^{(i,j)})+\sum_{u\in\mathcal{U}}R_{u}\hat{(\textbf{P}},\hat{\textbf{u}},\hat{\textbf{Z}}),\>$ (50)
@@
$\footnotesize\mathcal{P}_{5}:\max\sum_{d\in\mathcal{D}}R_{d}({\hat{\textbf{W}}},\textbf{u}^{(i,j)},\hat{\textbf{c}})+\sum_{u\in\mathcal{U}}R_{u}\hat{(\textbf{P}},\textbf{u}^{(i,j)},\hat{\textbf{Z}}).\>$ (52)
Although MOAN shares the similar structure like the other four networks, due to its small range, it is quite a challenge to keep them within the limited moving region while the meta-agent is performing the exploration. Hence, we have designed a customized regulator to ensure that u within the range of $[-A,A]$ where $A$ denotes the MA moving region, expressed as $\Delta\tilde{\textbf{u}}=\gamma\delta(\Delta\textbf{u})$ where $\gamma$ acts as an amplification operator and $\delta(.)$ represents the Tanh function. This design keeps the MA moving region within a restricted range during the inner iteration, allowing the updated region to be expressed as follows.
$\textbf{u}^{(i+1,j)}=\textbf{u}^{(i,j)}+\Delta\tilde{\textbf{u}}.\vspace{-0.2in}$ (53)
@@ IV-2 Outer iterations
$\footnotesize\mathcal{L}^{j}=-\mathcal{L}_{rate}^{j}+\mathcal{L}_{threshold}^{j}+\mathcal{L}_{norm}^{j}+\mathcal{L}_{common}^{j}+\mathcal{L}_{MA-dist}^{j},$ (54)
@@
$\footnotesize\mathcal{L}_{rate}^{j}=\sum_{d\in\mathcal{D}}R_{d}+\sum_{u\in\mathcal{U}}R_{u},\>$ (55)
@@
$\footnotesize\mathcal{L}_{threshold}^{j}=\sum_{d\in\mathcal{D}}\sum_{u\in\mathcal{U}}\rho_{1}\lambda_{1}(\mathcal{K}_{d}+\mathcal{K}_{u}),\>$ (56)
@@
where $\lambda_{2}(.)$ represents the indicator function and $\rho_{2}$ indicates the penalty weight. The fourth term of the meta-loss function replaces the constraints (30i) which can be expressed as follows,
@@
$\footnotesize\mathcal{A}_{n_{t}}=DS-||\textbf{t}_{n_{t}}^{BS}-\textbf{t}_{\hat{n_{t}}}^{BS}||\leq 0,1\leq n_{t}\neq\hat{n_{t}}\leq N_{T},$ (66)
$\footnotesize\mathcal{A}_{n_{r}}=DS-||\textbf{r}_{n_{r}}^{BS}-\textbf{r}_{\hat{n_{r}}}^{BS}||\leq 0,1\leq n_{t}\neq\hat{n_{t}}\leq N_{R},$ (67)
$\footnotesize{\lambda_{4}=}\begin{cases}{0},&\mbox{{if }}\mbox{{$\Omega$}},\\ {1},&\mbox{{if }}\mbox{{otherwise}},\end{cases}$ (68)
@@ IV-A Epoch iteration:
This iteration is responsible for updating the neural network parameters. After completing the designated number of outer iterations, the losses are accumulated and averaged accordingly, ¯$\footnotesize\bar{\mathcal{L}}=\frac{1}{N_{o}}\sum_{j=1}^{N_{o}}\mathcal{L}^{j}.$ Then, backward propagation is conducted and the Adam optimizer is used to update the neural networks embedded in these networks, as depicted below:
@@
Complexity Analysis: Following [19], the complexity of the proposed algorithm can be describes as follows. The complexity of the proposed GML algorithm lies in calculating the computations for the achievable rates and five neural networks. The first network, UTPON mainly relies on (10) and (11). Considering the dimension of $\textbf{Z}_{u,j}$, $\textbf{Z}_{U}$ is $N_{R}\times 1$ and hence the computational complexity of multiplying it with $\textbf{h}_{u}\in N_{R}\times 1$ is $\mathcal{O}(N_{R})$. This calculation is performed $12(U-1)+$ 1 times for computing the achievable rate and thus, the overall complexity becomes $\mathcal{O}(2(U-1)+1)^{2}N_{R})$. As gradient calculation in neural network occurs concurrently and thus, the computational complexity of the neural network is approximately $\mathcal{O}(2(U-1)+1)$. Hence, overall complexity of UTPON is $\mathcal{O}(2(U-1)+1)^{2}N_{R})$. Similarly, the computational complexity of BVN and CRN is $\mathcal{O}((D+1)^{2}N_{T})$, and RBN is $\mathcal{O}(2(U-1)+1)^{2}N_{R})$. Meanwhile, computational complexity of MOAN depends on the computation of (6), (7), (10), (11) and also its associated calculation of gradient. Hence, its computational complexity is $\mathcal{O}((D+1)^{2}N_{T})+\mathcal{O}(2(U-1)+1)^{2}N_{R})$. Finally, considering the inner, outer, and epoch iterations, the final time complexity of the proposed algorithm is $N_{e}N_{o}N_{i}(\mathcal{O}(2(U-1)+1)^{2}N_{R})+\mathcal{O}((D+1)^{2}N_{T})+\mathcal{O}((2(U-1)+1)^{2}N_{R})+\mathcal{O}((D+1)^{2}N_{T})+\mathcal{O}((2(U-1)+1)^{2}N_{R})).$
1 Randomly initialize $\theta_{\textbf{P}},\theta_{\textbf{W}},\theta_{\textbf{Z}},\theta_{\textbf{C}},\theta_{\textbf{u}},\textbf{P}^{(0,1)}$, $\textbf{W}^{(0,1)},\textbf{Z}^{(0,1)},\textbf{C}^{(0,1)}$, $\textbf{u}^{(0,1)}$;
2 Initialize inner iterations, $N_{i}$, outer iterations, $N_{o}$, and epoch iterations, $N_{e}$;
@@ 5       MAX$=0$;
7             $\textbf{P}^{(0,j)}=\textbf{P}^{(0,1)}$;
8            $\textbf{W}^{(0,j)}=\textbf{W}^{(0,1)}$ ;
9             $\textbf{Z}^{(0,j)}=\textbf{Z}^{(0,1)}$;
10             $\textbf{C}^{(0,j)}=\textbf{C}^{(0,1)}$;

### v3

@@ III Problem Formulation
$\text{s.t.}\qquad\textbf{t}_{n_{t}}^{BS}\in\mathcal{C}_{t}^{BS},\textbf{r}_{n_{r}}^{BS}\in\mathcal{C}_{r}^{BS},$ (31b)
@@
[…]
@@
where $\textbf{P}=[P_{u,j},P_{U}],\forall u\in\mathcal{U},\forall j\in J$, $\textbf{W}=[\textbf{W}_{c},\textbf{W}_{p,d}],\forall d\in\mathcal{D}$, $\textbf{Z}=[\textbf{Z}_{u,j},\textbf{Z}_{U}],\forall u\in\mathcal{U},\forall j\in J$, $\textbf{t}=[\textbf{t}_{n_{t}}^{BS},\textbf{t}^{UL}],\forall n_{t}\in N_{T}$, $\textbf{r}=[\textbf{r}_{n_{r}}^{BS},\textbf{r}^{DL}],\forall n_{r}\in N_{R}$. Here, (30b) refers to the power budget constraint of BS. (30c) and (30d) refer to the power budget constraint of UL UEs. Meanwhile, (30e) ensures the QoS constraints in terms of data rate requirements in both DL and UL UEs. (31b) and (30f) limit the ranges of MA movements. (30g) and (30h) associate with the receiving beamforming vectors at the BS. (30i) represents the constraint for successful decoding of common stream by all DL UEs. Finally, (31c) and (31d) ensure that minimum inter-MA distance $DS$ at the BS for practical implementation. It should be noted that $\mathcal{P}_{1},\mathcal{P}_{2},\mathcal{P}_{3}$ are highly nonconvex optimization problem. Specifically, the objective function, QoS constraints, receiving beamforming constraint, and minimum inter-MA distance constraints make the optimization problem highly intractable. Moreover, the optimization problem results in large-scale optimization problems with an increasing number of variables with increasing UEs. Hence, traditional optimization algorithms will suffer from high computational complexity to solve the proposed optimization problem. Thus, we propose a learning-based approach called gradient-based meta-learning, aka GML, to solve the problem with low complexity.
@@ IV Proposed Solution: Gradient-based Meta Learning (GML) For Resource Allocation
Conventional data-driven meta-learning approaches, such as model-agnostic meta-learning (MAML) [yuan2020transfer], rely heavily on extensive offline pretraining and frequent online adaptations. However, their high energy consumption and sensitivity to variations in data distribution make them impractical for dynamic, latency-sensitive environments. To address these challenges, we employ a model-driven GML framework, as proposed in [zhu2024robust], which eliminates the need for pretraining and focuses on trajectory optimization rather than optimizing individual variables. This approach has been demonstrated to be effective for non-convex optimization problems [xia2022metalearning]. The GML algorithm follows a three-layer nested cyclic optimization structure, consisting of epoch, outer, and inner iterations, which are further detailed below.
@@ IV-1 Inner iterations
Inner iteration is responsible for target variable optimization. We have utilized five separate neural networks to optimize our six target variables. The first neural network is called UE transmission power network (UTPON) to optimize the transmission power of UL UEs and hence optimize P. The second neural network is responsible for optimizing the beamforming vectors W at BS, and it is called the beamforming vector network (BVN). The third network is called the receiving beamforming network (RBN), which is responsible for optimizing Z. The fourth neural network is called a common rate network (CRN), which is responsible for optimizing the common rate portion in each DL UE. Finally, the last neural network is called the movable antenna network (MOAN). As the variables t and r both have similar characteristics and a purpose to serve, we use only MOAN to optimize them and $\textbf{u}=[\textbf{t},\textbf{r}]$. In each of these networks, optimization of these variables $\textbf{P},\textbf{W},\textbf{Z},\textbf{c},\textbf{u}$ is updated one after another in each inner iteration. In each network, optimization of the target variable starts from the initial value, while the other variables are inherited from their corresponding network. The target variables are optimized alternatively in their corresponding network until convergence. Hence, the update process in the $j$-th outer iteration is formulated as,
@@
where $\textbf{P}^{(i,j)},\textbf{W}^{(i,j)},\textbf{Z}^{(i,j)}$$\textbf{c}^{(i,j)}$$\textbf{u}^{(i,j)}$ represent $\textbf{P},\textbf{W},\textbf{Z},\textbf{c},\textbf{u}$ in the $i$-th inner iteration of the $j$-th outer iteration. Each of the networks above is iterated $N_{i}$ times in an inner iteration. As the formulated problem is highly non-convex with coupled optimization variables, the meta-learning structure allows us to solve the optimization problem in a tractable discrete manner. The details of each neural network are provided below.
@@
where $\hat{\textbf{W}},\hat{\textbf{u}},\hat{\textbf{c}},\hat{\textbf{Z}}$ represent either the initialized or updated parameters. The variable P is updated in two steps. First, we compute the sum rate based on (38) where we optimize only the target variable and keep the other variables from initializing or updating from the other network. This sum rate is basically the $\mathcal{L}_{rate}^{(i,j)}$ which is stated in (56). Then we calculate the initial loss function $\mathcal{L}^{(i,j)}$ based on (54) (this has been discussed in details in the section below). We calculate an initial gradient of $\textbf{P}^{(i,j)}$ with respect to the initial loss function we calculated. This initial gradient of $\textbf{P}^{(i,j)}$ is fed into the neural network. Then, neural network outputs another new gradient/step $\Delta\textbf{P}_{(i,j)}$, which is added to $\textbf{P}^{(i,j)}$ to update it for the next iteration. This update can be given as follows,
${P}_{u,j}^{(i+1,j)}={P}_{u,j}^{(i,j)}+\Delta{P}_{u,j}^{(i,j)},$ (39)
@@
$X_{1}=\frac{P_{u}^{max}}{\sum_{j\in J}P_{u,j}^{(i+1,j)}}P_{u,j}^{(i+1,j)},$ (43)
$X_{2}=\frac{P_{U}^{max}}{P_{U}^{(i+1,j)}}P_{U}^{(i+1,j)}.$ (44)
@@
[…]
@@
$\textbf{W}^{*}=\textbf{W}^{(i,j)}+\Delta\textbf{W}^{(i,j)}.$ (46)
@@
$\textbf{W}^{(i+1,j)}=\sqrt{\frac{P_{BS}}{\textrm{Tr}(\textbf{W}^{*}{\textbf{W}^{*}}^{H})}}\textbf{W}^{*}.$ (47)
Receiving Beamforming Network (RBN): In RBN, we optimize Z, while keeping other variables from their the initial states or updated from the other networks. Hence, the objective and update of gradient is measured as follows:
$\mathcal{P}_{6}:\max\sum_{d\in\mathcal{D}}R_{d}({\hat{\textbf{W}}},\hat{\textbf{u}},\hat{\textbf{c}})+\sum_{u\in\mathcal{U}}R_{u}\hat{(\textbf{P}},\hat{\textbf{u}},{\textbf{Z}^{(i,j)}}),\>$ (48)
$\textbf{Z}^{(i+1,j)}=\textbf{Z}^{(i,j)}+\Delta\textbf{Z}^{(i,j)}.$ (49)
@@
[…]
@@
[…]
Although MOAN shares the similar structure like the other four networks, due to its small range, it is quite a challenge to keep them within the limited moving region while the meta-agent is performing the exploration. Hence, we have designed a customized regulator to ensure that u within the range of [A,A][-A,A] where $A$ denotes the MA moving region, expressed as $\Delta\tilde{\textbf{u}}=\gamma\delta(\Delta\textbf{u})$ where $\gamma$ acts as an amplification operator and $\delta(.)$ represents the Tanh function. This design keeps the MA moving region within a restricted range during the inner iteration, allowing the updated region to be expressed as follows.
$\textbf{u}^{(i+1,j)}=\textbf{u}^{(i,j)}+\Delta\tilde{\textbf{u}}$. (53)
@@ IV-2 Outer iterations
$\mathcal{L}^{j}=-\mathcal{L}_{rate}^{j}+\mathcal{L}_{threshold}^{j}+\mathcal{L}_{norm}^{j}+\mathcal{L}_{common}^{j}+\mathcal{L}_{MA-dist}^{j},$ (54)
@@
$\mathcal{L}_{rate}^{j}=\sum_{d\in\mathcal{D}}R_{d}+\sum_{u\in\mathcal{U}}R_{u},\>$ (55)
@@
$\mathcal{L}_{threshold}^{j}=\sum_{d\in\mathcal{D}}\sum_{u\in\mathcal{U}}\rho_{1}\lambda_{1}(\mathcal{K}_{d}+\mathcal{K}_{u}),\>$ (56)
@@
where $\lambda_{2}(.)$ represents the indicator function and $\rho_{2}$ indicates the penalty weight. The fourth term of the meta-loss function replaces the constraints (30i) which can be expressed as,
@@
[…]
${\lambda_{4}=}\begin{cases}{0},&\mbox{{if }}\mbox{{$\Omega$}},\\ {1},&\mbox{{if }}\mbox{{otherwise}},\end{cases}$ (68)
@@ IV-A Epoch iteration:
This iteration is responsible for updating the neural network parameters. After completing the designated number of outer iterations, the losses are accumulated and averaged accordingly. $\bar{\mathcal{L}}=\frac{1}{N_{o}}\sum_{j=1}^{N_{o}}\mathcal{L}^{j}.$ Then, backward propagation is conducted, and the Adam optimizer is used to update the neural networks embedded in these networks, as depicted below:
[…]
@@
Complexity Analysis: Following [zhu2024robust], the complexity of the proposed algorithm can be described as follows. The complexity of the proposed GML algorithm lies in calculating the computations for the achievable rates and five neural networks. The first network, UTPON mainly relies on (10) and (11). Considering the dimension of $\textbf{Z}_{u,j}$, $\textbf{Z}_{U}$ is $N_{R}\times 1$ and hence the computational complexity of multiplying it with $\textbf{h}_{u}\in N_{R}\times 1$ is $\mathcal{O}(N_{R})$. This calculation is performed 2(U1)+12(U-1)+1 times for computing the achievable rate and thus, the overall complexity becomes $\mathcal{O}(2(U-1)+1)^{2}N_{R})$. As gradient calculation in neural network occurs concurrently, the computational complexity of the neural network is approximately $\mathcal{O}(2(U-1)+1)$. Hence, the overall complexity of UTPON is $\mathcal{O}(2(U-1)+1)^{2}N_{R})$. Similarly, the computational complexity of BVN and CRN is $\mathcal{O}((D+1)^{2}N_{T})$, and RBN is $\mathcal{O}(2(U-1)+1)^{2}N_{R})$. Meanwhile, the computational complexity of MOAN depends on the computation of (6), (7), (10), (11) and also its associated calculation of gradient. Hence, its computational complexity is $\mathcal{O}((D+1)^{2}N_{T})+\mathcal{O}(2(U-1)+1)^{2}N_{R})$. Finally, considering the inner, outer, and epoch iterations, the final time complexity of the proposed algorithm is $N_{e}N_{o}N_{i}(\mathcal{O}(2(U-1)+1)^{2}N_{R})+\mathcal{O}((D+1)^{2}N_{T})+\mathcal{O}((2(U-1)+1)^{2}N_{R})+\mathcal{O}((D+1)^{2}N_{T})+\mathcal{O}((2(U-1)+1)^{2}N_{R})).$
1Initialize $\theta_{\textbf{P}},\theta_{\textbf{W}},\theta_{\textbf{Z}},\theta_{\textbf{C}},\theta_{\textbf{u}},\textbf{P}^{(0,1)}$, $\textbf{W}^{(0,1)},\textbf{Z}^{(0,1)},\textbf{C}^{(0,1)}$, $\textbf{u}^{(0,1)}$, inner iterations, $N_{i}$, outer iterations, $N_{o}$, epoch iterations, $N_{e}$;
@@ 3 $\bar{\mathcal{L}}=0$,MAX=0=0;
5    $\textbf{P}^{(0,j)}=\textbf{P}^{(0,1)}$$\textbf{W}^{(0,j)}=\textbf{W}^{(0,1)}$ ;
6    $\textbf{Z}^{(0,j)}=\textbf{Z}^{(0,1)}$, $\textbf{C}^{(0,j)}=\textbf{C}^{(0,1)}$;